# Impact of reflection Comptonization on X-ray reflection spectroscopy: the case of EXO 1846–031

### v1

@@ Abstract
Within the disk-corona model, it is natural to expect that a fraction of reflection photons from the disk are Compton scattered by the hot corona (reflection Comptonization), even if this effect is usually ignored in X-ray reflection spectroscopy studies. We study the effect by using NICER and NuSTAR data of the Galactic black hole EXO 1846–031 in the hard-intermediate state with the model simplcutx. Our analysis suggests that a scattered fraction of order 10% is required to fit the data, but the inclusion of reflection Comptonization does not change appreciably the measurements of key-parameters like the black hole spin and the inclination angle of the disk.
@@ I Introduction
EXO 1846–031 is a black hole X-ray binary discovered by EXOSAT on April 3, 1985 (Parmar and White, 1985). After a quiescent period of over two decades, it experienced a new outburst in 2019, which was first detected by MAXI/GSC on July 23, 2019 Negoro et al. (2019). This outburst was also observed by other missions during the following few weeks, such as Swift Mereminskiy et al. (2019), NICER (Bult et al., 2019), and NuSTAR Miller et al. (2019). During the hard to soft state transition (Yang et al., 2019a), the source showed a spectrum with strong reflection features Miller et al. (2019) and quasi-periodic oscillations (QPOs) (Yang et al., 2019b). Draghis et al. (2020) found that the black hole in EXO 1846–031 has a spin parameter very close to 1 and the viewing angle of its accretion disk is high, close to 70 deg. Thanks to the very strong reflection features in the NuSTAR spectrum, this source has been widely used to test new reflection models (see, e.g., Refs. Abdikamalov et al. (2021); Tripathi et al. (2021a)) as well as different spacetime metrics from modified gravity theories (see, e.g., Refs. Tripathi et al. (2021b); Yu et al. (2021); Gu et al. (2022); Tao et al. (2023)). Abdikamalov et al. (2021) showed that a radial disk ionization profile can improve the fit of the spectrum around 7 keV. QPOs were studied in Liu et al. (2021) and the possible existence of a disk wind was studied in Wang et al. (2021).
@@ II Observations and data reduction
Figure 2: Light curve of EXO 1846–031 from Monitor of All-sky X-ray Image (MAXI) [^2] Light curve data:
@@
. The red-dashed line marks the observation date of the data we used.
@@ III Data Analysis
We use XSPEC v12.12.1 (Arnaud, 1996) to analyze the spectra. We use two different datasets for the spectral fitting process: the spectra from NuSTAR only (our “NuSTAR data”), and the spectra from both NuSTAR and NICER (our “NuSTAR+NICER data”) for a joint fitting. The reason is to demonstrate the influence of soft X-ray data on the fitting. The fits are made across the 3.0-79.0 keV energy band of NuSTAR and the 1.0-10.0 keV energy band for NICER. In the joint fitting of NuSTAR and NICER data, NuSTAR data below 4.5 keV are ignored in order to avoid the mismatch between NICER and NuSTAR at the low energy band, as demonstrated by other studies (Tao et al., 2019; Nath et al., 2024).
Firstly, we fit the data with an absorbed disk black-body and power-law to see the reflection features in the spectrum. In XSPEC language, the model reads
@@
where tbabs describes the galactic absorption with the abundance table from Wilms et al. (2000), diskbb (Pringle, 1981; Mitsuda et al., 1984) describes the thermal component, and nthcomp (Zdziarski et al., 1996; Życki et al., 1999) describes the power-law component. We link $T_{\rm in}$ in diskbb to $T_{\rm bb}$ in nthcomp. We clearly see reflection features with an a broad iron line around 7 keV and a Compton hump peaking around 20-30 keV (Fig. 3).
Figure 3: Data to best-fit model ratio plots for an absorbed disk blackbody and power-law model. The left panel is for NuSTAR data and the right panel is for NuSTAR+NICER data. The blue and purple crosses are for the FPMA/NuSTAR and FPMB/NuSTAR data and the green crosses are for the NICER data.
@@ III.1 Model 1: conventional modeling
Previous study on the same observation shows that a reflection model that assumes a radial ionization gradient can fit the data better (Abdikamalov et al., 2021). Therefore, we implement the relxillion_nk model which describes the ionization profile with a power-law:
@@ III.2 Model 2: self-consistent modeling with simplcutx
Adopting a more self-consistent model structure, the power-law component is generated from Comptonization of the thermal component in the corona, and the reflection component originating from the disk surface will also be Comptonized (Steiner et al., 2017). In XSPEC language, the model expression is tbabs*simplcutx*(diskbb+relxillionCp_nk) (Model 2). We set the reflection fraction $R_{\rm F}$ in simplcutx as 1 to avoid parameter degeneracy. In simplcutx, $R_{\rm F}$ is defined as the number of scattered photons that return to illuminate the disk divided by the number of photons that reach infinity.
@@ III.4 Other setting and final results
For the NuSTAR data, we only set a cross-calibration constant between FPMA and FPMB. For the NuSTAR+NICER data, we add an edge component to fit an absorption edge around 1.84 keV. We also apply the xscat model (Smith et al., 2016) for all data as the field of view of NICER is comparable to the typical size of a dust scattering halo. We choose the MRN dust model for the scattering. For the extraction regions, the radius is ${}^{\prime\prime}$ for the NuSTAR data and ${}^{\prime\prime}$ for the NICER data, in accordance with the configuration in Barillier et al. (2023). Furthermore, we apply a cross-calibration model called crabber (Steiner et al., 2010) which is designed to standardize the calibration parameters between different detectors for the Crab Nebula (power-law normalization and index) via a multiplicative power-law correction. All parameters in the NuSTAR and NICER spectra are linked except the cross-calibration parameters.
@@
Table 3: Best-fit values of all models with NuSTAR data. $P$ means that we reach the lower or upper limit of the model.
@@
Table 4: Best-fit values and unabsorbed flux of all models with NuSTAR+NICER data. $P$ means that we reach the lower or upper limit of the model. In the estimate of the Eddington ratio (last row), $D$ is the distance in the unit of kpc and $M$ is the mass in the unit of $M_{\odot}$. The estimate of the total flux is obtained by integrating the total spectrum of the best-fit model from 0.01 to 1000 keV.
@@ IV.1 Impact of reflection Comptonization
Figure 7: The 1-D $\chi^{2}$ contours for black hole spin and inclination angle for Model 1A and Model 2A with NuSTAR data. The horizontal dash line represents the $1\sigma$, $2\sigma$, $3\sigma$ confidence levels for a single parameter.
@@ IV.2 Impact of soft X-ray data
From the comparison of the results from the analyses of the NuSTAR and NuSTAR+NICER data, we see the impact of the soft X-ray band in the analysis of the data. The disk component peaks around 2.0-3.0 keV (see Fig. 6), but NuSTAR lacks coverage below 3.0 keV. Consequently, the fit of the disk component with NuSTAR data may not be as accurate as the fit with NuSTAR+NICER data. In terms of parameters estimation, while some parameters differ slightly (e.g. $N_{\rm H}$, $T_{\rm in}$), key parameters such as spin and inclination angle still do not change significantly. It is tentatively suggested that the inclusion of soft X-ray data does not significantly impact the parameter estimations of the best-fit model. However, by including NICER soft X-ray data, we identify an additional $\chi^{2}$ minimum, not as deep, which exhibits a more reasonable flux of the disk component despite having a slightly higher AICc; see the columns of Models 1A-2 and 2A-2 in Table 4.
The mass and distance of EXO 1846-031 have not been dynamically measured yet. Studies of spectral analysis suggest mass measurements ranging from 3.24 to 7.1-12.43 $M_{\odot}$ (Strohmayer and Nicer Observatory Science Working Group, 2020; Nath et al., 2024), and distance measurements ranging from 2.4 to 7 kpc (Parmar et al., 1993; Williams et al., 2022). Considering a distance of 7 kpc and a mass in the range 5 to 12 $M_{\odot}$, the Eddington ratio $L/L_{\rm Edd}$ is 21.6%–51.7% for Model 1A-1, 18%–43% for Model 2A-1, 7.7%–13% for Model 1A-2, and 7.8%–13% for Model 2A-2. Furthermore, the unabsorbed disk component is much fainter in Models 1A-2 and 2A-2 (refer to the “disk/total” raw in Table 4). This indicates that for certain values of mass and distance, the luminosity of Models 1A-1 and 2A-1 is too high, while the results of Models 1A-2 and 2A-2 are more reasonable.
@@
where $\kappa\sim 1.7$ is the ratio of color temperature to effective temperature (Shimura and Takahara, 1995), then we can obtain $R_{\rm in}$ which is the true inner radius. For a Kerr black hole, $R_{\rm ISCO}$ (radius of the innermost stable circular orbit) can be derived from the spin (Bardeen et al., 1972). Since we apply X-ray reflection spectroscopy to measure the spin, we should identify the $R_{\rm in}$ with $R_{\rm ISCO}$ (Bambi et al., 2021). Although the estimates of $R_{\rm in}$ and $R_{\rm ISCO}$ from observed data have significant uncertainty due to the lack of accurate measurements of mass and distance, $R_{\rm in}$ should not be much larger than $R_{\rm ISCO}$ given the reflection-fitting results. $R_{\rm in}$ inferred from the results of Model 1A-1 and 2A-1 is about 9.7-70 $R_{\rm ISCO}$ (with the same uncertainty of mass and distance in estimation of Eddington ratio), while from the results of Model 1A-2 and 2A-2 is about 1.0-2.7 $R_{\rm ISCO}$. Therefore, the best fits of Model 1A-2 and 2A-2 are somewhat favored.
@@ V Conclusions
1. Considering reflection Comptonization by using simplcutx does not yield a significantly different best fit and it does not affect the values of key parameters such as black hole spin and inclination angle. In the case of coronal properties evolving during state transitions, the variation in $f_{\rm SC}$ potentially providing a useful consistency check.
2. The inclusion of soft X-ray data does not significantly affect the estimations of key parameters, but gives a new fit with more reasonable luminosity, inner radius of disk and emissivity profile. The fit of NuSTAR+NICER data with a weak disk component and non-zero $q_{\rm out}$ is the most reasonable in this case.
Acknowledgments – This work was supported by the Natural Science Foundation of Shanghai, Grant No. 22ZR1403400, and the National Natural Science Foundation of China (NSFC), Grant No. 12250610185, 11973019, and 12261131497.

### v2

@@ Abstract
Within the disk-corona model, it is natural to expect that a fraction of reflection photons from the disk are Compton scattered by the hot corona (reflection Comptonization), even if this effect is usually ignored in X-ray reflection spectroscopy studies. We study this effect by using NICER and NuSTAR data of the Galactic black hole EXO 1846–031 in the hard-intermediate state with the model simplcutx. Our analysis suggests that a scattered fraction of order 10% is required to fit the data, but the inclusion of reflection Comptonization does not change appreciably the measurements of key-parameters like the black hole spin and the inclination angle of the disk.
@@ I Introduction
EXO 1846–031 is a black hole X-ray binary discovered by EXOSAT on April 3, 1985 (Parmar and White, 1985). After a quiescent period of over two decades, it experienced a new outburst in 2019, which was first detected by MAXI/GSC on July 23, 2019 Negoro et al. (2019). This outburst was also observed by other missions during the following few weeks, such as Swift Mereminskiy et al. (2019), NICER (Bult et al., 2019), and NuSTAR Miller et al. (2019). During the hard to soft state transition (Yang et al., 2019a), the source showed a spectrum with strong reflection features Miller et al. (2019) and quasi-periodic oscillations (QPOs) (Yang et al., 2019b). Draghis et al. (2020) found that the black hole in EXO 1846–031 has a spin parameter very close to 1 and the inclination angle of its accretion disk (i.e., the angle between the normal to the disk surface and our line of sight) is high, close to 70 deg. Thanks to the very strong reflection features in the NuSTAR spectrum, this source has been widely used to test new reflection models (see, e.g., Refs. Abdikamalov et al. (2021); Tripathi et al. (2021a)) as well as different spacetime metrics from modified gravity theories (see, e.g., Refs. Tripathi et al. (2021b); Yu et al. (2021); Gu et al. (2022); Tao et al. (2023)). Abdikamalov et al. (2021) showed that a radial disk ionization profile can improve the fit of the spectrum around 7 keV. QPOs were studied in Liu et al. (2021) and the possible existence of a disk wind was studied in Wang et al. (2021).
@@ II Observations and data reduction
Figure 2: Light curve of EXO 1846–031 from Monitor of All-sky X-ray Image (MAXI). The red-dashed line marks the observation date of the data we used. Light curve data available from:
@@ III Data Analysis
We use XSPEC v12.12.1 (Arnaud, 1996) to analyze the spectra. We use two different datasets for the spectral fitting process: the spectra from NuSTAR only (our “NuSTAR data”), and the spectra from both NuSTAR and NICER (our “NuSTAR+NICER data”) for a joint fitting. The reason is to demonstrate the influence of soft X-ray data on the fitting. The fits are made across the 3.0-79.0 keV energy band of NuSTAR and the 1.0-10.0 keV energy band for NICER. In the joint fitting of NuSTAR and NICER data, NuSTAR data below 4.5 keV are ignored in order to avoid the mismatch between NICER and NuSTAR in the low energy band. Indeed, if we fit the data and we plot the residuals, we clearly see that the NICER and NuSTAR spectra are not consistent. Even in the study presented in Ref. Nath et al. (2024) on these NICER and NuSTAR spectra, the authors ignore the NuSTAR data below 4 keV and a similar choice is adopted in Ref. Tao et al. (2019), where the authors analyze NuSTAR and Swift data of GRS 1716–249 and ignore the NuSTAR data below 4.5 keV. We do not know the exact origin of such a discrepancy in the low energy band; it may be some calibration issue in NuSTAR Madsen et al. (2020); Chakraborty et al. (2021).
First, we fit the data with an absorbed disk black-body and power-law to see the reflection features in the spectrum. In XSPEC language, the model reads
@@
where tbabs describes the galactic absorption with the abundance table from Wilms et al. (2000), diskbb (Pringle, 1981; Mitsuda et al., 1984) describes the thermal component, and nthcomp (Zdziarski et al., 1996; Życki et al., 1999) describes the power-law component. We link $T_{\rm in}$ in diskbb to $T_{\rm bb}$ in nthcomp. We clearly see reflection features with a broad iron line around 7 keV and a Compton hump peaking around 20-30 keV (Fig. 3).
Figure 3: Data to best-fit model ratio plots for an absorbed disk blackbody and power-law model. The left panel is for the NuSTAR data and the right panel is for the NuSTAR+NICER data. The blue and purple crosses are for the FPMA/NuSTAR and FPMB/NuSTAR data and the green crosses are for the NICER data.
@@ III.1 Model 1: conventional modeling
Previous study of the same observation shows that a reflection model that assumes a radial ionization gradient can fit the data better (Abdikamalov et al., 2021). Therefore, we implement the relxillion_nk model that describes the ionization profile with a power-law:
@@ III.2 Model 2: self-consistent modeling with simplcutx
Adopting a more self-consistent model structure, the power-law component is generated from Comptonization of the thermal component in the corona, and the reflection component originating from the disk surface will also be Comptonized (Steiner et al., 2017). In XSPEC language, the model expression is tbabs*simplcutx*(diskbb+relxillionCp_nk) (Model 2). We set the reflection fraction $R_{\rm F}$ in simplcutx to 1 to avoid parameter degeneracy with the scattered fraction $f_{\rm SC}$aaaThe scattering fraction $f_{\rm SC}$ is the fraction of thermal disk photons that are Compton-scattered by the corona into the power-law component. The reflection fraction $R_{\rm F}$ describes the flux of Compton-scattered photons directed back to the disk relative to the Compton flux that is transmitted to infinity. These two parameters are normally degenerate. See Ref. Steiner et al. (2017) for more details..
@@ III.4 Other setting and final results
For the NuSTAR data, we only set a cross-calibration constant between FPMA and FPMB. For the NuSTAR+NICER data, we add an edge component to fit an absorption edge around 1.84 keV. We also apply the xscat model (Smith et al., 2016) for all data as the field of view of NICER is comparable to the typical size of a dust scattering halo. We choose the MRN dust model for the scattering. For the extraction regions, the radius is 50′′ for the NuSTAR data and 180′′ for the NICER data, in accordance with the configuration in Barillier et al. (2023). Furthermore, we apply a cross-calibration model called crabber (Steiner et al., 2010) which is designed to standardize the calibration parameters between different detectors for the Crab Nebula (power-law normalization and index) via a multiplicative power-law correction. All parameters in the NuSTAR and NICER spectra are linked except the cross-calibration parameters.
@@
Table 3: Best-fit values of all models with NuSTAR data. $P$ means that we reach the lower or upper limit of the model. The reported uncertainties correspond to the 90% confidence intervals.
@@
Table 4: Best-fit values and unabsorbed fluxes of all models with NuSTAR+NICER data. All models have $q_{\rm out}$ free in the fit and there is no distant reflector (case A). Models 1A-1 and 2A-1 are the fits with the lowest values of $\chi^{2}$ (respectively without and with simplcutx). Models 1A-2 and 2A-2 are local minima, as described in Section IV.2 (respectively without and with simplcutx). $P$ means that we reach the lower or upper limit of the model. In the estimate of the Eddington ratio (last row), $D$ is the distance in kpc and $M$ is the mass in units of $M_{\odot}$. The estimate of the total flux is obtained by integrating the total spectrum of the best-fit model from 0.01 to 1000 keV. The reported uncertainties correspond to the 90% confidence intervals.
@@ IV.1 Impact of reflection Comptonization
Figure 7: The 1-D $\chi^{2}$ contours for black hole spin (left) and inclination angle (right) for Model 1A and Model 2A with NuSTAR data. The horizontal dashed line represents the $1\sigma$, $2\sigma$, $3\sigma$ confidence levels for a single parameter.
@@ IV.2 Impact of soft X-ray data
From the comparison of the results from the analyses of the NuSTAR and NuSTAR+NICER data, we see the impact of the soft X-ray band in the analysis of the data. The disk component peaks around 2.0-3.0 keV (see Fig. 6), but NuSTAR lacks coverage below 3.0 keV. Consequently, the fit of the disk component with NuSTAR data may not be as accurate as the fit with NuSTAR+NICER data. In terms of parameter estimation, while some parameters differ slightly (e.g. $N_{\rm H}$, $T_{\rm in}$), key parameters such as spin and inclination angle still do not change significantly. It is tentatively suggested that the inclusion of soft X-ray data does not significantly impact the parameter estimations of the best-fit model. However, by including NICER soft X-ray data, we identify an additional $\chi^{2}$ minimum, not as deep, which exhibits a more appropriate flux of the disk component despite having a slightly higher AICc; see the columns of Models 1A-2 and 2A-2 in Table 4. The main difference between 1A-1/2A-1 and 1A-2/2A-2 is in the flux of the thermal component of the disk. In the observations analyzed in this work, the source was in a hard-intermediate state. In 1A-1/2A-1, the Norm parameter of diskbb exceeds $10^{5}$ and this component counts for about 50% of the total unabsorbed flux. In Wang et al. (2021), the authors report the analysis of EXO 1846–031 in the soft state and find a value for the Norm parameter around 1500. These results are difficult to explain, as we should expect that the thermal component is stronger in the soft state. Moreover, from a very high value of the Norm parameter, we would infer that the inner radius is far from the radius of the innermost stable circular orbit (ISCO), see Eq. (5) below, which is the opposite of what we infer from the analysis of the reflection features. If we believe that the correct fits are 1A-2/2A-2, Norm is about 1000 and makes sense: the disk component is not stronger than in the soft state and the inner edge of the disk is around the ISCO.
The mass and distance of EXO 1846–031 have not been dynamically measured yet. Studies of spectral analysis suggest mass measurements ranging from 3.24 to 7.1-12.43 $M_{\odot}$ (Strohmayer and Nicer Observatory Science Working Group, 2020; Nath et al., 2024), and distance measurements ranging from 2.4 to 7 kpc (Parmar et al., 1993; Williams et al., 2022). Considering a distance of 7 kpc and a mass in the range 5 to 12 $M_{\odot}$, the Eddington ratio $L/L_{\rm Edd}$ is 21.6%–51.7% for Model 1A-1, 18%–43% for Model 2A-1, 7.7%–13% for Model 1A-2, and 7.8%–13% for Model 2A-2. Furthermore, the unabsorbed disk component is much fainter in Models 1A-2 and 2A-2 (refer to the “disk/total” row in Table 4). This indicates that for certain values of mass and distance, the luminosity of Models 1A-1 and 2A-1 is too high, while the results of Models 1A-2 and 2A-2 are more appropriate.
@@
where $\kappa\sim 1.7$ is the ratio of color temperature to effective temperature (Shimura and Takahara, 1995), then we can obtain $R_{\rm in}$ which is the true inner radius. For a Kerr black hole, $R_{\rm ISCO}$ can be derived from the spin (Bardeen et al., 1972). Since we apply X-ray reflection spectroscopy to measure the spin, we should identify the $R_{\rm in}$ with $R_{\rm ISCO}$ (Bambi et al., 2021). Although the estimates of $R_{\rm in}$ and $R_{\rm ISCO}$ from observed data have significant uncertainty due to the lack of accurate measurements of mass and distance, $R_{\rm in}$ should not be much larger than $R_{\rm ISCO}$ given the reflection-fitting results. $R_{\rm in}$ inferred from the results of Model 1A-1 and 2A-1 is about 9.7-70 $R_{\rm ISCO}$ (with the same uncertainty of mass and distance in estimation of Eddington ratio), while from the results of Model 1A-2 and 2A-2 is about 1.0-2.7 $R_{\rm ISCO}$. Therefore, the best fits of Model 1A-2 and 2A-2 are somewhat favored.
@@ V Conclusions
1. Considering reflection Comptonization by using simplcutx does not yield a significantly different best fit and it does not affect the values of key parameters such as black hole spin and inclination angle. In the case of coronal properties evolving during state transitions, the variation in $f_{\rm SC}$ potentially provides a useful consistency check.
2. The inclusion of soft X-ray data does not significantly affect the estimations of key parameters, but gives a new fit with more appropriate luminosity, inner radius of disk and emissivity profile. The fit of NuSTAR+NICER data with a weak disk component and non-zero $q_{\rm out}$ is the most appropriate in this case.
Acknowledgments – This work was supported by the Natural Science Foundation of Shanghai, Grant No. 22ZR1403400, and the National Natural Science Foundation of China (NSFC), Grants No. 12250610185, No. 11973019, and No. 12261131497.